Pcl5 reversabile reaction pcl3+ cl2
At 300 k kc  is 0.204 mol/l what is the value of k

Grade:11

1 Answers

Vikas TU
3 years ago
For a gaseous reaction 2B→A, the equilibrium constant Kp is less than Kc.
Kp = Kc (RT)^del n 
del n = 1-2 = -1 
Kp = Kc(RT)^-1 
Kp is less than Kc 
Kp = 0.204/300 * 0.083 = 5.644*10^-5